OpenStudy (anonymous):

3x^3-12x^2-17x+10= 0

OpenStudy (anonymous):

i assume you want the zeros of this using factor theorem one root is x=5

OpenStudy (anonymous):

u can find the others by dividing x-5 into your expression to get a quadratic - solve this to get the other 2 roots

OpenStudy (anonymous):

\[3x^3-12x^2-17x+10 = 0\]\[(x-5)(3x^2+3x-2) = 0\] so \[x = 5\space or \space \frac{-3 \pm \sqrt{3^2-4(3)(-2}}{6}\] \[x = 5 \space or \space \frac{1}{6}(-3 \pm \sqrt{33})\]